OpenStudy (anonymous):

Solve for P= ab+2/c for b

OpenStudy (radar):

ab=P-2/c b=(P-2c)/a

OpenStudy (anonymous):

ab=P-2/c Multiply both sides by c (to get rid of the denominator) cab = cP-2 divide both sides by ac b = (cP-2)/ac